the direction says to "solve each equation.check your answer.if appropiate, write identity or no solution."

2 (4 - 2r) = -2(r +5)

*could you please show the steps also*

Answer by roxy9229(32)

*=times

2(4-2r) = -2(r+5)
distributive property
distribute the 2 on the left
& the -2 on the right
2*4=8 2*2r=4r -2*r=-2r -2*5=-10
change the +(-10) into - 10 to make it simpler
8-4r = -2r-10
+4r +4r
8 = 2r-10
add 10 from both sides
18=2r
divide both sides by 2
9=r
or r=2